# Quantized topological transport mediated by the long-range couplings

### v1

@@ Abstract
Topological physics uncovers promising approaches to control the localization and propagation of light by tailoring the bandstructure of the material and harnessing localized or propagating topological modes [1, 2, 3].
Even richer physics arises when the Hamiltonian of the system varies in time. If such variation is periodic, this gives rise to Floquet physics and non-equilibrium phases with tailored properties [4, 5]. Of special interest is a periodic variation of the Hamiltonian resulting in the transport of excitations analogously to the water flow driven by the Archimedean screw. Leveraging the topological nature of the system such transport can be made disorder-resilient in the sense that the charge transferred during a single pumping cycle is robustly quantized [6, 7].
Historically the first example of such topological transport was the so-called Thouless pump [8] which utilized the Rice-Mele model [6, 7] with time-varying on-site energies and couplings between the sites. The latter system, in turn, is a generalization of the celebrated Su-Schrieffer-Heeger model [9], a paradigmatic example of a one-dimensional topological system.
The Thouless pump has been realized experimentally for ultracold atoms in a dynamically controlled optical lattice [10, 11, 12] as well as for photonic systems [13, 14, 15, 16]. Recent generalizations include systems with strong nonlinearity [17, 18, 19] and pumping of multiband systems [20, 21, 22] which could feature non-Abelian physics [20, 21]. Importantly, the protocol of Thouless pumping in simple two-band systems requires a synchronized variation of both couplings and on-site energies, which is quite challenging to implement in the arrays of evanescently coupled optical waveguides requiring a simultaneous modulation of the refractive index contrast and distance between the waveguides [16].
@@
where the couplings $J_{1,2}$ and $t_{1,2}$ depend on time according to Eqs. (2)-(5).
The spectrum for a finite 14-unit-cell lattice in Fig. 2(a) provides an intuition on how the system behaves in the adiabatic limit, i.e. when the driving frequency is much smaller than the characteristic eigenfrequencies of the system. The results suggest that the gap in the spectrum remains open throughout the entire pumping cycle, and there are only two edge states which cross the bandgap and are responsible for the quantized transport. In Figure 2(a) the energies of left- and right-localized edge states are shown by red and blue colors, respectively.
In addition to the instantaneous spectrum we also examine the instantaneous eigenvectors. While Bloch functions do not provide immediate insights into the properties of the pumping scheme, their linear combinations known as Wannier functions can be readily visualized; they feature good localization, form an orthogonal set and are related to each other via translation [7].
To construct the Wannier functions, we first introduce the position operator in the form suitable for the periodic systems [23, 24, 25]
@@
Figure 2: Instantaneous spectrum and Wannier centers of the driven system. (a) Evolution of the instantaneous spectrum calculated for a finite system consisting of $N=14$ unit cells. Left- and right-localized edge states are highlighted by red and blue, respectively. (b) Evolution of the instantaneous Wannier centers for the same system during a single driving cycle.
@@
The eigenvectors of the projected position operator $\hat{X}_{p}$ provide the Wannier functions of the lowest band [7]. At the same time, the eigenvalues of $\hat{X}_{p}$, $\lambda_{m}$, define so-called Wannier centers via
@@
We plot the instantaneous Wannier centers for our system in Fig. 2(b) and observe that all Wannier centers shift by one unit cell during the pumping cycle. Since an arbitrary input state from the lowest band can be decomposed into the superposition of Wannier functions, this suggests that the designed protocol transports any such state by one unit cell during a single pumping cycle.
@@
So far we discussed the adiabatic limit of our model. To assess how the transport protocol works at a finite driving frequency, we solve the temporal evolution of the system consisting of $N=14$ unit cells. We choose the driving period $T=7$, which substantially exceeds the average inverse bandgap size $1/\Delta\simeq 0.6$, justifying the adiabatic approximation. We assume that the initial state is localized in the middle of the array and belongs entirely to the lowest band, i.e.
@@
Solving the Schrödinger equation, we recover the probability distribution in Fig. 3(a) which shows that initially localized wavepacket starts to spread over the lattice. While the distribution $|\psi_{n}(t)|^{2}$ itself does not exhibit clear signatures of transport, it is instructive to track the motion of the wavepacket center of mass, i.e. $\left<x(t)\right>=\left<\psi\left|\hat{x}\right|\psi\right$> [Fig. 3(b)]. We observe that the center-of-mass motion resembles the temporal dependence of the Wannier centers and exhibits a shift equal to 1.005 unit cell during a single pumping cycle. Nearly quantized change of $\left<x\right$> is due to the topological nature of the pump, while slight violation of the quantization appears due to the non-adiabatic evolution of the system. Note that in order to have strictly quantized pumping at a finite driving frequency the temporal profile of the driving has to be fine-tuned [26] using, for instance, counter-adiabatic driving method [27, 28].
@@
Figure 4: (a) Chosen geometry of the lattice and its spatial modulation. Insets show electric field profiles (amplitude) of symmetric and antisymmetric modes in vertical and detuned horizontal waveguides, respectively. (b) Modulation of the couplings $J_{1,2}$, $t_{1,2}$ and detunings $u_{1,2}$ extracted from numerically simulated splitting of the dimer eigenmodes. (c) Instantaneous spectrum of the propagation constants computed for a finite lattice with $14$ unit cells. Left- and right-localized edge states are highlighted by red and blue, respectively. (d) Tight-binding simulation of the real-space discrete diffraction pattern for the initial point-like excitation projected onto the lowest Bloch band of a $14$ unit cell lattice and modulation period $L=50$ cm. (e) Time dependence of the center of mass of the intensity distribution (black solid line), displaced by $1.0068$ unit cells over one modulation period. The trajectory of the Wannier center is shown by the dashed red curve for comparison.
@@
As a specific platform to implement our protocol of topological pumping at optical wavelengths, we propose an optical waveguide lattice fabricated via the femtosecond laser writing technique [29, 30, 31]. We exploit a formal analogy between the tight-binding Schrödinger equation and the coupled-mode description of the waveguide lattice: the time variable $t=\tau T$ in the former corresponds to the propagation distance $z=\tau L$ in the waveguide system, where $L$ is the spatial period of the modulation along the waveguides.
@@
We propose the zigzag-like geometry shown in Fig. 4(a) consisting of the main vertical waveguides shown in white and detuned connector waveguides depicted in red. The positioning and orientation of connector waveguides is chosen in such a way that they efficiently couple to the sublattice of main waveguides close to them, but remain practically decoupled from another sublattice of main waveguides. Depending on parameters, the waveguides can support the modes with the different symmetry of the near field profile. While horizontally oriented connector waveguides could utilize symmetric ($s$) mode, antisymmetric ($p_{x}$) modes odd under $\sigma_{v}$ reflection generally enable larger bandgap (see Supplementary Materials). Accordingly, we choose the refractive index contrast of connector waveguides such that the $s$-like modes of the vertical waveguides have the propagation constant closer to the $p$-like modes of the connector sites, thus realizing an instance of multi-orbital physics [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Alternatively, for the connector $p$-mode sites one could use fine-tuned photonic molecules based on pairs of closely placed vertically oriented waveguides [36].
Since the connector waveguides are assumed to be detuned from the main ones, they realize non-resonant coupling and can be excluded from the description using the degenerate perturbation theory [48]. By doing so, we recover an effective Hamiltonian possessing next-nearest-neighbor couplings $t_{1,2}=\kappa+\gamma^{2}/(\kappa+\Delta)$ between the main waveguides belonging to the same sublattice. In addition, main waveguides also acquire additional detunings $u_{1,2}=-\gamma^{2}/(\kappa+\Delta)$, assuming $|\Delta|\gg\gamma$. Here, $\Delta=k_{z}^{p}-k_{z}^{s}$ is the detuning in the propagation constants of $p$-modes in horizontal connector waveguides ($k_{z}^{p}$) from vertical $s$-modes in main lattice waveguides ($k_{z}^{s}$), $\gamma$ is the absolute value of the coupling between main and detuned waveguides, and $\kappa$ is the direct coupling between the waveguides belonging to the same sublattice. Importantly, the antisymmetric shape of $p$-like modes ensures that the couplings $t_{1,2}$ and detunings $u_{1,2}$ are modulated out of phase. As a result, the two mechanisms of bandgap opening add up, lead to the larger bandgap and allow to reduce the modulation period.
@@
Such bending of the waveguides gives rise to the $z$-dependent couplings depicted in Fig. 4(b). Here, the effective couplings $J_{1,2}$, $t_{1,2}$, $\kappa$ and detunings $u_{1,2}$ are recovered by computing the eigenmodes of a pair of waveguides and tracking the splitting between the modes, which quantifies the strength of the coupling. The obtained values of the couplings justify the use of the degenerate perturbation theory for the chosen detuning $\Delta$.
@@
Next we simulate numerically the spectrum for a lattice consisting of $14$ unit cells [Fig. 4(c)] and observe a good agreement with the tight-binding result using numerically calculated couplings (see Supplementary Materials for details). The spectrum shows left-and right-localized topological edge states traversing the complete bandgap in the opposite directions, indicating the presence of the quantized pump. Moreover, the spectrum in this case features a particularly large bandgap, facilitating quantized transport for smaller lattices and shorter modulation periods with the greater robustness to the disorder.
@@
Finally, the tight-binding simulation of topological pumping in Fig. 4(d) shows the asymmetric spreading of the initial point-like excitation projected onto the lower Bloch band for the array consisting of $14$ unit cells with the modulation period $L=50$ cm. The center-of-mass trajectory of the intensity distribution in Fig. 4(e) (black curve) shows quantized transport and follows quite closely the trajectory of the respective Wannier center (red dashed curve) with the total shift of $1.0068$ unit cells per modulation period. For smaller modulation periods, quantized nature of the transport is violated resulting, for instance, in the shift of $0.84$ unit cell for the modulation period $L=20$ cm. Note that the lattice periods required for quantized adiabatic transport may be further decreased for lattices operating close to the flat-band or all-bands-flat conditions [34], and also in the nonlinear pumping regime when the diffraction of the wave packet is suppressed by the nonlinearity [17].
@@
In summary, we have proposed a topological pump allowing to transfer the quantum state by varying nearest-neighbor and next-nearest-neighbor couplings. This protocol provides an alternative to the celebrated Thouless pump and does not involve the change of on-site energies, which is especially suitable for the arrays of evanescently coupled optical waveguides. The proposed scheme is not limited to optical waveguides and can be readily generalized to the other physical platforms, e.g. to the polariton condensates trapped in a slowly varying periodic potential [49, 50] with a conceptual scheme elaborated in the Supplementary Materials.
@@
See Supplementary Materials for the details of numerical simulations of the system with symmetric and anti-symmetric connector modes in an optical waveguide lattice, including lattice geometry and chosen modulation scheme, application of the degenerate perturbation theory for connector sites, evaluation of the effective couplings and detuning of connector waveguides, comparison of the tight-binding and numerical finite-lattice spectra and tight-binding simulations of quantized pumping as well the proposed implementation for the lattice of coupled polariton condensates.
@@
We acknowledge Andrei Stepanenko and Anton Nalitov for valuable discussions. Theoretical models were supported by the Russian Science Foundation (Grant No. 24-72-10069). Numerical simulations were supported by Priority 2030 Federal Academic Leadership Program.

### v2

@@ Abstract
Topological physics uncovers promising approaches to control the localization and propagation of light by tailoring the bandstructure of the material and harnessing localized or propagating topological modes [Lu2014, Ozawa2019, Price2022].
Even richer physics arises when the Hamiltonian of the system varies in time. If such variation is periodic, this gives rise to Floquet physics and non-equilibrium phases with tailored properties [Goldman2014, Rechtsman2013]. Of special interest is a periodic variation of the Hamiltonian resulting in the transport of excitations analogously to the water flow driven by the Archimedean screw. Leveraging the topological nature of the system such transport can be made disorder-resilient in the sense that the charge transferred during a single pumping cycle is robustly quantized [BookShen, Asboth].
Historically the first example of such topological transport was the so-called Thouless pump [Thouless1983] which utilized the Rice-Mele model [BookShen, Asboth] with time-varying on-site energies and couplings between the sites. The latter system, in turn, is a generalization of the celebrated Su-Schrieffer-Heeger model [Su1979Jun], a paradigmatic example of a one-dimensional topological system.
The Thouless pump has been realized experimentally for ultracold atoms in a dynamically controlled optical lattice [Lohse2015, Nakajima2016, Nakajima2021] as well as for photonic systems [Kraus2012, Tang2016, Fedorova2020, Cerjan2020]. Recent generalizations include systems with strong nonlinearity [Sebabrata2021, Fu2022, Fu2022a, Sebabrata2023, Ravets2025], fractional Thouless pumping [Tao2025a], pumping of two-dimensional [Wang2022] and multiband systems [You2022, Sun2024, Tao2025] which could feature non-Abelian physics [You2022, Sun2024]. Importantly, the protocol of Thouless pumping in simple two-band systems requires a synchronized variation of both couplings and on-site energies, which is quite challenging to implement in the arrays of evanescently coupled optical waveguides requiring a simultaneous modulation of the refractive index contrast and distance between the waveguides [Cerjan2020].
@@
which features two bulk bands separated by the bandgap, while the couplings $J_{1,2}$ and $t_{1,2}$ depend on time according to Eqs. (2)-(5).
The spectrum for a finite 14-unit-cell lattice with the open boundary conditions in Fig. 2(a) provides an intuition on how the system behaves in the adiabatic limit, i.e. when the driving frequency is much smaller than the characteristic eigenfrequencies of the system. The results suggest that the gap in the spectrum remains open throughout the entire pumping cycle, and there are only two edge states which cross the bandgap and exhibit nonzero group velocity. As we discuss in the Supplementary Materials, these edge states are a signature of the topological tranport.
A complementary perspective is obtained by studying the same lattice but with the periodic boundary conditions. While Bloch functions do not provide immediate insights into the properties of the pumping scheme, their linear combinations known as Wannier functions can be readily visualized; they feature good localization, form an orthogonal set and are related to each other via translation [Asboth].
Figure 2: Instantaneous spectrum and Wannier centers of the driven system. (a) Evolution of the instantaneous spectrum calculated for a finite system consisting of $N=14$ unit cells with open boundary conditions. Left- and right-localized edge states are highlighted by red and blue, respectively. (b) Evolution of the instantaneous Wannier centers for the same system with periodic boundary conditions
during a single driving cycle.
To construct the Wannier functions, we first introduce the position operator in the form suitable for the periodic systems [Resta1998Mar, Marzari1997Nov, Marzari2012Oct]
@@
The eigenvectors of the projected position operator $\hat{X}_{p}$ provide the Wannier functions of the lowest band [Asboth]. At the same time, the eigenvalues of $\hat{X}_{p}$, $\lambda_{m}$, define so-called Wannier centers via
@@
We plot the instantaneous Wannier centers for our system in Fig. 2(b) and observe that all Wannier centers shift by one unit cell during the pumping cycle. Since an arbitrary input state from the lowest band can be decomposed into the superposition of Wannier functions, this suggests that the designed protocol transports any such state by one unit cell during a single pumping cycle. The topological nature of this pumping scheme can be further highlighted by computing the space-time Chern number of the lowest band which is $C=1$ (see Supplementary Materials).
@@
So far we discussed the adiabatic limit of our model. To assess how the transport protocol works at a finite driving frequency, we solve the temporal evolution of the system consisting of $N=14$ unit cells with the open boundary conditions. We choose the driving period $T=7$, which substantially exceeds the average inverse bandgap size $1/\delta\simeq 0.6$, justifying the adiabatic approximation. We assume that the initial state is localized in the middle of the array and belongs entirely to the lowest band, i.e.
@@
Solving the Schrödinger equation, we recover the probability distribution in Fig. 3(a) which shows that initially localized wavepacket starts to spread over the lattice. While the distribution $|\psi_{n}(t)|^{2}$ itself does not exhibit clear signatures of transport, it is instructive to track the motion of the wavepacket center of mass, i.e. $\left<x(t)\right>=\left<\psi\left|\hat{x}\right|\psi\right$> [Fig. 3(b)]. We observe that the center-of-mass motion resembles the temporal dependence of the Wannier centers and exhibits a shift equal to 1.005 unit cell during a single pumping cycle. Nearly quantized change of $\left<x\right$> is due to the topological nature of the pump, while slight violation of the quantization appears due to the non-adiabatic evolution of the system. Note that in order to have strictly quantized pumping at a finite driving frequency the temporal profile of the driving has to be fine-tuned [Malikis2022] using, for instance, counter-adiabatic driving method [Angelis2020, Ke2025]. Due to the chosen timespan, the wavepacket initialized in the middle of the lattice practically does not interact with the edges and thus open boundary conditions do not destroy quantized transport.
@@
Figure 4: (a) Chosen geometry of the lattice and its spatial modulation. Insets show electric field profiles (amplitude) of symmetric and antisymmetric modes in vertical and detuned horizontal waveguides, respectively. Insets illustrate the choice of the unit cell with identical spacings and with $p$-mode photonic molecule connectors (a2), the adiabatic modulation of waveguides along their propagation coordinate $z$ for one unit cell (a3) and a 55-unit-cell lattice (a4). (b) Modulation of the couplings $J_{1,2}$, $t_{1,2}$ and detunings $u_{1,2}$ extracted from numerically simulated splitting of the dimer eigenmodes. (c) Instantaneous spectrum of the propagation constants computed for a finite lattice with $14$ unit cells. Left- and right-localized edge states are highlighted by red and blue, respectively. (d) Tight-binding simulation of the real-space discrete diffraction pattern for the initial point-like excitation projected onto the lowest Bloch band of a $14$ unit cell lattice and modulation period $L=50$ cm. (e) Time dependence of the center of mass of the intensity distribution (black solid line), displaced by $1.0068$ unit cells over one modulation period. The trajectory of the Wannier center is shown by the dashed red curve for comparison.
@@
As a specific platform to implement our protocol of topological pumping at optical wavelengths, we propose an optical waveguide lattice fabricated via femtosecond laser writing technique [Szameit2010Jul, Wang2021Mar, Yan2024Oct]. We exploit a formal analogy between the tight-binding Schrödinger equation and the coupled-mode description of the waveguide lattice: the time variable $t=\tau T$ in the former corresponds to the propagation distance $z=\tau L$ in the waveguide system, where $L$ is the spatial period of the modulation along the waveguides.
@@
We propose the zigzag-like geometry shown in Fig. 4(a) consisting of the main vertical waveguides shown in white and detuned connector waveguides depicted in red. The positioning and orientation of connector waveguides is chosen in such a way that they efficiently couple to the sublattice of main waveguides close to them, but remain practically decoupled from another sublattice of main waveguides. Depending on parameters, the waveguides can support the modes with the different symmetry of the near field profile. To enable larger bandgap, we require that the connector waveguides support dipolar $p_{x}$ modes spectrally close to the $s$ modes of the main waveguides (see Supplementary Materials). This realizes an instance of multi-orbital physics [Vicencio2025Jul, Guzman-Silva2021Aug, Caceres-Aravena2022Jun, Caceres-Aravena2024Sep, Mazanov2024Apr, Caceres-Aravena2019Jul, Caceres-Aravena2020Aug, Savelev2020Oct, Mikhin2023Mar, Gorbach2023Apr, Liu2024Dec, Gao2024Mar, Rajeevan2025Jan, Mazanov2024May, Li2016Sep, Pelegri2019Feb]. As the fabrication of elliptical waveguides with the different ellipse orientations is challenging, for the connector $p$-mode sites one could use fine-tuned photonic molecules based on pairs of closely placed vertically oriented waveguides [Mazanov2024Apr].
Since the connector waveguides are assumed to be detuned from the main ones, they realize non-resonant coupling and can be excluded from the description using the degenerate perturbation theory [Bir1974, Gorlach2018Mar, Mittal2019Oct, Mukherjee2018Aug] We illustrate this procedure for a trimer of two $s$-mode main waveguides connected by the detuned $p$-mode waveguide, described by the Hamiltonian
$H_{\text{trimer}}=\begin{pmatrix}0&\kappa&\gamma\\ \kappa&0&-\gamma\\ \gamma&-\gamma&\Delta\end{pmatrix}$, (12)
where $\gamma$ is the absolute value of the coupling between main and detuned waveguides, $\kappa$ is the direct coupling between the two main waveguides, and $\Delta=k_{z}^{p}-k_{z}^{s}$ is the propagation constants detuning. Assuming $|\Delta|\gg\kappa,\gamma$, we exclude the detuned $p$-mode [Bir1974, Gorlach2018Mar] and obtain the effective $2\times 2$ effective Hamiltonian
$H_{\text{dimer}}=\begin{pmatrix}u&t\\ t&u\end{pmatrix}$ (13)
with effective next-nearest-neighbor couplings $t=\kappa+\gamma^{2}/(\kappa+\Delta)$ and detunings $u=-\gamma^{2}/(\kappa+\Delta)$, as further discussed in the Supplementary Materials.
Importantly, the antisymmetric shape of $p$-like modes ensures that the couplings $t_{1,2}$ and detunings $u_{1,2}$ for each sublattice are modulated out of phase. As a result, the two mechanisms of bandgap opening add up, leading to the larger bandgap and allowing to reduce the modulation period.
@@
Such bending of the waveguides gives rise to the $z$-dependent couplings as further analyzed in the Supplementary Materials. To calculate these dependencies, we first obtain the couplings $J(dx)$, $\gamma(dy)$, $\kappa$ at perfect degeneracy ($\Delta=0$) as functions of respective waveguide spatial displacements from the eigenmodes of respective waveguide pairs, tracking the splitting between the modes which quantifies the strength of the couplings. Substituting $\gamma(dy)$, $\kappa$ and $\Delta$ into the effective couplings $t_{1,2}$ and detunings $u_{1,2}$ and using specific modulations $dx(\tau),dy(\tau)$, we arrive to the results in Fig. 4(b).
@@
Next we simulate numerically the spectrum for a lattice consisting of $14$ unit cells [Fig. 4(c)] and observe a good agreement with the tight-binding result using numerically calculated couplings (see Supplementary Materials for details). The spectrum shows left-and right-localized topological edge states traversing the complete bandgap in the opposite directions, indicating the presence of topological pump. Moreover, the spectrum in this case features a particularly large bandgap, facilitating quantized transport for smaller lattices and shorter modulation periods with the greater robustness to the disorder.
@@
Finally, the tight-binding simulation of topological pumping in Fig. 4(d) shows the asymmetric spreading of the initial point-like excitation projected onto the lower Bloch band for the array consisting of $14$ unit cells with the modulation period $L=50$ cm. The center-of-mass trajectory of the intensity distribution in Fig. 4(e) (black curve) shows quantized transport and follows quite closely the trajectory of the respective Wannier center (red dashed curve) with the total shift of $1.0068$ unit cells per modulation period. Figures 4(d,e) feature many parallels with Fig. 3(a), even though the former describes a realistic model with the detunings and non-harmonic coupling dependencies on $\tau$, while the latter – a simple model with harmonic coupling modulation without detunings. This can be attributed to the adiabatic regime of the pumping with similar adiabaticity parameters $\alpha=\delta L$, where $\delta$ is the bandgap size, as well as the validity of the degenerate perturbation theory.
For smaller modulation periods, quantized nature of the transport is violated resulting, for instance, in the shift of $0.84$ unit cell for the modulation period $L=20$ cm. Note that the lattice periods required for quantized adiabatic transport may be further decreased for lattices operating close to the flat-band or all-bands-flat conditions [Caceres-Aravena2022Jun], and also in the nonlinear pumping regime when the diffraction of the wave packet is suppressed by the nonlinearity [Sebabrata2021].
@@
In summary, we have proposed a topological pump allowing to transfer the quantum state by varying nearest-neighbor and next-nearest-neighbor couplings. This protocol provides an alternative to the celebrated Thouless pump and does not involve the change of on-site energies, which is especially suitable for the arrays of evanescently coupled optical waveguides. The proposed scheme is not limited to optical waveguides and can be readily generalized to the other physical platforms, e.g. to the polariton condensates trapped in a slowly varying periodic potential [Gnusov2023, Fraser2024] with a conceptual scheme elaborated in the Supplementary Materials.
@@
See Supplementary Materials for the details of numerical simulations of the system with symmetric and anti-symmetric connector modes in an optical waveguide lattice, including lattice geometry and chosen modulation scheme, application of the degenerate perturbation theory for connector sites, evaluation of the effective couplings and detuning of connector waveguides, comparison of the tight-binding and numerical finite-lattice spectra and tight-binding simulations of quantized pumping, proposal for photonic molecule connector waveguides, results for the full tight-binding model, as well the proposed implementation for the lattice of coupled polariton condensates.
@@
We acknowledge Andrei Stepanenko and Anton Nalitov for valuable discussions. This work was supported by the Russian Science Foundation grant No. 25-79-31027.